OpenStudy (anonymous):

simplify the expression -3(b-7)

OpenStudy (luigi0210):

Distribute each term: \[\LARGE -3(b-7)=-3(b)-3(-7)=?\]

OpenStudy (texaschic101):

do you understand Luigi's explanation ? Are you still having trouble ?

OpenStudy (texaschic101):

whats -3 * b ?

OpenStudy (anonymous):

b - 10?

OpenStudy (texaschic101):

no... -3 * b = -3b Now whats -3 * -7 ? Do not add, you are multiplying. Remember, when multiplying by two negatives, it changes to a positive. So what is -3 * -7 ?

OpenStudy (anonymous):

right 21

OpenStudy (texaschic101):

so now we have : -3b + 21 understand ?

OpenStudy (anonymous):

thanks yes

OpenStudy (texaschic101):

just multiply the number outside of the parenthesis with everything inside the parenthesis.

OpenStudy (texaschic101):

doing this will get rid of the parenthesis

OpenStudy (texaschic101):

you got anymore ?